# Should I Share this Translation? Evaluating Quality Feedback for User Reliance on Machine Translation

### v2

@@ 4.2 RQ2: Which Feedback is Most Effective?
4.3 RQ3: Which Feedback do Users Rely on Most?
We compute switch percentage to capture participants’ behavioral patterns of reliance in Figure 5. Under-reliance is highest in the error highlights group (25.6%), followed by backtranslation (21.1%), LLM explanation (18.1%), and QA table (15.3%). Interestingly, implicit feedback types (QA table (7.78%) and backtranslation (7.22%)) yield lower over-reliance than explicit ones (LLM explanation (13.3%) and error highlights (10.0%)). Across all conditions, participants are more likely to maintain their initial decisions (regardless of correctness) than to change them, as under-reliance consistently exceeds over-reliance, and appropriate reliance (no switch) exceeds (switch).
We further examine switch percentages by shareability label in Appendix G.2. Participants show higher over-reliance and lower under-reliance for shareable examples than non-shareable ones. [^11] We refer to shareable as examples labeled “Safe to share as-is,” and non-shareable as those labeled “Needs bilingual review before sharing”. This suggests that participants are more likely to change their decisions when initially judging translations as shareable, but tend to maintain their decisions when judging them as non-shareable.
@@ 5 Analysis
As shown in Figure 4, participants consistently achieve significantly higher decision accuracy (p < 0.01 for error highlights; p < 0.001 for others) and CWA scores (p < 0.001) on shareable examples than on non-shareable ones, except in the LLM explanation group. This suggests that participants make more accurate and appropriate decisions when evaluating good translations than problematic ones, indicating that helping users reliably identify critical MT errors remains a challenge.
@@
In the post-task survey, participants rated the quality feedback they received in terms of perceived mental burden, helpfulness, and trust for future use. As shown in Table 1, QA table group reported the highest level of trust (4.22), aligning with findings from Section 4.2 that this feedback is most effective at improving appropriate reliance. In contrast, participants who received error highlights gave the lowest ratings for helpfulness (3.83) and trust (3.89), which is consistent with their relatively poor performance in both decision accuracy and appropriate reliance (CWA). Interestingly, while LLM explanations do not yield large gains in decision accuracy or CWA, they have relatively high ratings for helpfulness (4.39) and trust (4.11), which may reflect the over-reliance discussed in Section 4.3. In terms of mental burden, the error highlights group reported higher score (4.94) than other groups. We attribute this to the nature of error highlights as a target-side feedback mechanism (Leiter et al., 2024), which displays highlights on the Spanish MT output, making it difficult to interpret for monolingual source speakers.
Feedback Mental burden (↓) Helpfulness (↑) Trust (↑)
[Uncaptioned image] 4.94 3.83 3.89
[Uncaptioned image] 4.06 4.39 4.11
[Uncaptioned image] 4.06 4.39 4.06
[Uncaptioned image] 4.00 4.39 4.22
Таблица 1: Average mental burden (1-7), helpfulness (1-5), and trust for future use (1-5) for each quality feedback type. Best scores for each metric are bold.
@@ 5.3 What Makes Quality Feedback Helpful?
We present participants’ responses on how they used quality feedback in their decision-making and which aspects they found helpful or unhelpful. For error highlights, two participants found them largely unhelpful, noting that the highlights were only shown on the Spanish MT, illustrating a key limitation of a target-side feedback. However, some appreciated the explicitness of the highlights, stating that they pointed to ‘‘areas that not have been accurately translated’’ or ‘‘key translation mistakes’’. Similarly, four participants valued LLM explanations for showing ‘‘exactly what is correct or incorrect’’. Some further appreciated for offering insights into better phrasing, alternative translations, or contextual relevance.
In contrast, participants who received implicit feedback described a more self-directed decision-making process. Backtranslation was considered helpful for verifying ‘‘some tiny unsure details’’ or checking whether ‘‘the core meaning of the original English text was preserved in the Spanish MT’’ by comparing the two English texts. Similarly, QA table encouraged participants to revisit the MT output when a mismatch was detected (‘‘if the statement is blue, I double check the phrase again’’), ‘‘compare the words and make determinations’’ themselves. Detailed comments are provided in Appendix G.4.
@@ 6 Conclusion
We explore the utility of quality feedback in helping monolingual source speakers make reliable MT decisions. We conduct a human study where participants decide whether Spanish MT outputs are safe to share, first independently and then with one of four feedback types on MT quality: two explicit (error highlights, LLM explanation) and two implicit (backtranslation, QA table). We find that all feedback types except error highlights significantly improve decision accuracy and appropriate reliance (§4.1). Implicit feedback, especially QA table, outperforms explicit feedback in both objective performance and self-reported ratings (§4.2, §5.2) and while explicit feedback prompts more decision changes, it also increases over-reliance (§4.3). We further show that participants are better at confirming good translations than detecting problematic ones (§5.1).
Overall, our findings underscore the value of feedback that guides users’ implicit interpretation rather than prescribing decisions. Implicit methods may be especially effective, as they preserve users’ agency in the decision-making process (Savoldi et al., 2025). Along with insights into what participants found helpful across feedback types (§5.3), our work calls for future efforts to explore optimal feedback designs and combinations that help users reliably identify critical MT errors.
7 Limitation
@@ Acknowledgements
  • Bo et al. (2025) Jessica Y Bo, Sophia Wan, and Ashton Anderson. 2025. To rely or not to rely? evaluating interventions for appropriate reliance on large language models. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems, CHI ’25, New York, NY, USA. Association for Computing Machinery.
@@
  • Guo et al. (2024) Ziyang Guo, Yifan Wu, Jason Hartline, and Jessica Hullman. 2024. A decision theoretic framework for measuring AI reliance. pages 221–236.
  • Guo et al. (2025a) Ziyang Guo, Yifan Wu, Jason Hartline, and Jessica Hullman. 2025a. Unexploited information value in human-AI collaboration. Preprint, arXiv:2411.10463.
  • Guo et al. (2025b) Ziyang Guo, Yifan Wu, Jason Hartline, and Jessica Hullman. 2025b. The value of information in human-AI decision-making. Preprint, arXiv:2502.06152.
@@
  • Pitman (2021) Jeff Pitman. 2021. Google translate: One billion installs, one billion stories. https://blog.google/products/translate/new-features-make-translate-more-accessible-for-its-1-billion-users/. Engineering Manager, Google Translate.
@@
  • Schemmer et al. (2022) Max Schemmer, Patrick Hemmer, Niklas Kühl, Carina Benz, and Gerhard Satzger. 2022. Should i follow AI-based advice? measuring appropriate reliance in human-AI decision-making. Preprint, arXiv:2204.06916.
@@
  • Somers (2005) Harold Somers. 2005. Round-trip translation: What is it good for? In Proceedings of the Australasian Language Technology Workshop 2005, pages 127–133.
  • Srinivasan and Thomason (2025) Tejas Srinivasan and Jesse Thomason. 2025. Adjust for trust: Mitigating trust-induced inappropriate reliance on AI assistance. Preprint, arXiv:2502.13321.
@@ C.1 Annotation Interface
Post-Task Survey.
In the post-task survey, we dynamically replace feedback with the participant’s assigned quality feedback type. The survey includes the following three questions:
@@ C.2 Recruitment Process
We conducted our user studies on the Prolific platform. For the pre-screening survey, we used the same language proficiency questions as in the pre-task survey (Appendix C) to reliably recruit monolingual English speakers. To ensure high-quality responses, we limited participation to users with a Prolific approval rate above 90% and at least 10 prior submissions. We recruited 165 participants and invited 98 who selected ‘‘I can read long and difficult texts in English without help’’ for English proficiency, and either ‘‘I cannot understand any Spanish words or sentences at all’’ or ‘‘I can read some Spanish words and very simple sentences, such as greetings and common expressions’’ for Spanish proficiency. Each pre-screening participant received 0.20 USD, totaling 44.53 USD including Prolific platform fees.
@@
Of the 98 invited participants, 73 completed the main task. Each received a base payment of 5 USD for 20 minutes of participation, with an additional 2 USD performance-based bonus for those achieving over 70% overall decision accuracy. A total of 25 participants qualified for this bonus. One participant who failed both attention check questions was compensated but excluded from the final analysis, resulting in 72 valid responses. The median task completion time was 15 minutes, corresponding to an effective pay rate of 20 USD per hour. The total cost for the main task, including Prolific fees, was 486.67 USD.
@@
Приложение F Potential Learning Effects
While our original two-step setup (Independent → AI-Assisted) follows a widely adopted setup in prior work on AI-assisted decision making (Guo et al., 2024, 2025a, 2025b), particularly in studies focused on measuring reliance and trust in AI systems (Buçinca et al., 2021; Schemmer et al., 2022; Mehandru et al., 2023; Bo et al., 2025; Srinivasan and Thomason, 2025), we conducted an additional human study to directly address the concern of potential practice effects. In this true repetition control condition (Independent → Independent), participants completed the same task without receiving any quality feedback between steps. Specifically, we conducted (1) a pre-screening survey to recruit participants matching the original user demographics (English as first/primary/fluent language, low or no Spanish fluency, and residing in the U.S.), and (2) the main task, where each participant viewed the same 20 examples twice in succession and made shareability and confidence judgments both times without any feedback in between.
We recruited 18 participants for this control condition, all of whom self-reported English as their first language, with an average English proficiency of 5/5 and an average Spanish proficiency of 1.83/5. The task took a median of 12 minutes and 30 seconds to complete.
As shown in Table 3, both decision accuracy and confidence-weighted accuracy (CWA) in Step 2 for the true control group are significantly lower than those in all other feedback conditions (M = 0.605, S.E. = 0.036 for accuracy; M = 0.172, S.E. = 0.054 for CWA). Additionally, the control group’s performance in Step 1 did not differ significantly from that of the other condition groups. These findings suggest that the gains observed in the feedback conditions are not attributed to repeated exposure to the same example, but rather to the quality feedback provided in Step 2. Moreover, the within-group difference between Step 1 and Step 2 in the true control group showed no statistically significance, indicating that simply seeing the same example again without any feedback does not lead to better decision-making nor appropriate reliance.
In the post-survey, we asked the true control group participants to rate their experience on two dimensions: (1) Mental workload: ‘‘How much mental burden (e.g., thinking, deciding) did you experience while doing the task?’’, and (2) Helpfulness: ‘‘How helpful was seeing the same example twice in your Spanish translation quality assessment?’’ Compared to the feedback conditions, the true control group reported higher mental burden (M = 5.83) and lower perceived helpfulness (M = 2.89), suggesting that repeated exposure without any quality feedback increased cognitive load without enhancing perceived utility.
Group Condition Step 1 Acc. (↑) Step 2 Acc. (↑) Step 1 CWA (↑) Step 2 CWA (↑) Metal burden (1-7, ↓) Helpfulness (1-5, ↑)
True control - 0.613 (0.036) 0.605 (0.037) 0.182 (0.055) 0.172 (0.054) 5.83 2.89
Condition Error Highlights 0.593 (0.022) 0.647 (0.025) 0.156 (0.041) 0.239 (0.042) 4.94 3.83
LLM Explanation 0.634 (0.023) 0.693 (0.024) 0.232 (0.042) 0.353 (0.042) 4.06 4.39
Backtranslation 0.603 (0.024) 0.716 (0.025) 0.188 (0.041) 0.388 (0.043) 4.06 4.39
QA Table 0.659 (0.024) 0.765 (0.022) 0.282 (0.040) 0.488 (0.040) 4.00 4.39
Таблица 3: Comparison of the true control group and two-step groups across quality feedback types. Reported are mean values (with standard errors) for decision accuracy and CWA, as well as average ratings for two post-task survey questions: mental burden and helpfulness.
Приложение G Detailed Results
G.1 Independent vs. AI-Assisted
@@ G.2 Switch Percentage
Таблица 4: Categorization of decision transitions from Independent (rows) to AI-Assisted (columns). Appropriate: Appropriate reliance; Over: Over-reliance; Under: Under-reliance.
@@
Таблица 5: Detailed switch percentage results by quality feedback type and reliance (appropriate, over-, under-reliance).
G.3 Per Shareability Label
@@ G.4 Participant Responses
Таблица 6: Detailed results for decision accuracy and CWA by quality feedback type. Ind.: Independent decision-making step; AI: AI-assisted decision-making step. Values represent mean scores with corresponding 95% confidence intervals. Best scores for each column is bold.
@@
Таблица 7: Detailed results for decision accuracy and CWA by quality feedback type and shareability label. [Uncaptioned image]: Safe to share as-is; [Uncaptioned image]: Needs bilingual review before sharing. Values represent mean scores with corresponding 95% confidence intervals. Best scores for each column is bold.
@@
Таблица 8: Participants’ responses on how they used quality feedback in their decision-making process, specifically why the feedback was perceived as confusing or unhelpful ([Uncaptioned image]) or why it was considered helpful ([Uncaptioned image]).

### v3

@@ 4.2 RQ2: Which Feedback is Most Effective?
4.3 RQ3: Which Feedback do Users Rely on Most Appropriately?
For each quality feedback, we compute switch percentage to capture participants’ behavioral patterns of reliance in Figure 5. Under-reliance is highest in the error highlights group (25.6%), followed by backtranslation (21.1%), LLM explanation (18.1%), and QA table (15.3%). Interestingly, implicit feedback types (QA table (7.78%) and backtranslation (7.22%)) yield lower over-reliance than explicit ones (LLM explanation (13.3%) and error highlights (10.0%)). Across all conditions, participants are more likely to maintain their initial decisions (regardless of correctness) than to change them, as under-reliance consistently exceeds over-reliance, and appropriate reliance (no switch) exceeds (switch).
We further examine switch percentages by shareability label in Appendix F.2. Participants show higher over-reliance and lower under-reliance for shareable examples than non-shareable ones. [^11] We refer to shareable as examples labeled “Safe to share as-is,” and non-shareable as those labeled “Needs bilingual review before sharing”. This suggests that participants are more likely to change their decisions when initially judging translations as shareable, but tend to maintain their decisions when judging them as non-shareable.
@@ 5 Analysis
As shown in Figure 4, treatment condition participants consistently achieve significantly higher decision accuracy (p < 0.01 for error highlights; p < 0.001 for others) and CWA scores (p < 0.001) on shareable examples than on non-shareable ones. One exception is the LLM explanation group, where the differences are statistically not significant. This suggests that participants generally make more accurate and appropriate decisions when evaluating good translations than problematic ones, indicating that helping users reliably identify critical MT errors remains a challenge.
Condition Mental burden Helpfulness Trust
Control 5.83 2.89 -
Error Highlights 4.94 3.83 3.89
Explanation 4.06 4.39 4.11
Backtranslation 4.06 4.39 4.06
QA Table 4.00 4.39 4.22
Таблица 1: Average mental burden (1-7, ↓), helpfulness (1-5, ↑), and trust for future use (1-5, ↑) for each condition group. Best scores for each metric are bold.
@@
In the post-task survey, participants rated the quality feedback they received in terms of perceived mental burden, helpfulness, and trust for future use. As shown in Table 1, the control condition group reported highest mental burden (5.83) and lowest helpfulness (2.89), suggested that repeated exposure without any quality feedback increased cognitive load without enhancing perceived utility.
Among the four treatment conditions, QA table group reported the highest level of trust (4.22), aligning with findings from Section 4.2 that this feedback is most effective at improving appropriate reliance. In contrast, participants who received error highlights gave the lowest ratings for helpfulness (3.83) and trust (3.89), which is consistent with their relatively poor performance in both decision accuracy and appropriate reliance (CWA). Interestingly, while LLM explanations do not yield large gains in decision accuracy or CWA, they have relatively high ratings for helpfulness (4.39) and trust (4.11), which may reflect the over-reliance discussed in Section 4.3. In terms of mental burden, the error highlights group reported higher score (4.94) than other groups. We attribute this to the nature of error highlights as a target-side feedback mechanism (Leiter et al., 2024), which displays highlights on the Spanish MT output, making it difficult to interpret for monolingual source speakers.
@@ 5.3 What Makes Quality Feedback Helpful?
We present treatment condition participants’ responses on how they used quality feedback in their decision-making and which aspects they found helpful or unhelpful. For error highlights, two participants found them largely unhelpful, noting that the highlights were only shown on the Spanish MT, illustrating a key limitation of a target-side feedback. However, some appreciated the explicitness of the highlights, stating that they pointed to ‘‘areas that not have been accurately translated’’ or ‘‘key translation mistakes’’. Similarly, four participants valued LLM explanations for showing ‘‘exactly what is correct or incorrect’’. Some further appreciated for offering insights into alternative translations or contextual relevance.
In contrast, participants who received implicit feedback described a more self-directed decision-making process. Backtranslation was considered helpful for verifying ‘‘some tiny unsure details’’ or checking whether ‘‘the core meaning of the original English text was preserved in the Spanish MT’’ by comparing the two English texts. Similarly, QA table encouraged participants to revisit the MT output when a mismatch was detected (‘‘if the statement is blue, I double check the phrase again’’), ‘‘compare the words and make determinations’’ themselves. Detailed comments are provided in Appendix F.4.
@@ 6 Conclusion
We explore the utility of quality feedback in helping monolingual source speakers make reliable MT decisions. We conduct a between-subjects human study where participants decide whether Spanish MT outputs are safe to share, first independently and then with one of five conditions. The four treatment conditions include different types of MT quality feedback: two explicit (error highlights and LLM explanation) and two implicit (backtranslation and QA table).
We find that all feedback types except error highlights significantly improve decision accuracy and appropriate reliance (§4.1). Implicit feedback, especially QA table, outperforms explicit feedback in both objective performance and self-reported ratings (§4.2, §5.2) and while explicit feedback prompts more decision changes, it also increases over-reliance (§4.3). We further show that participants are better at confirming good translations than detecting problematic ones (§5.1).
Overall, our findings underscore the value of feedback that guides users’ implicit interpretation rather than prescribing decisions. Implicit methods may be especially effective, as they preserve users’ agency in the decision-making process (Savoldi et al., 2025; Xiao et al., 2025a). Along with insights into what participants found helpful across feedback types (§5.3), our work calls for further research exploring feedback to help users reliably identify critical MT errors in realistic use cases.
7 Limitations
@@ Acknowledgements
  • Agrawal et al. (2022) Sweta Agrawal, Nikita Mehandru, Niloufar Salehi, and Marine Carpuat. 2022. Quality estimation via backtranslation at the WMT 2022 quality estimation task. In Proceedings of the Seventh Conference on Machine Translation (WMT), pages 593–596, Abu Dhabi, United Arab Emirates (Hybrid). Association for Computational Linguistics.
@@
  • Briakou et al. (2023) Eleftheria Briakou, Navita Goyal, and Marine Carpuat. 2023. Explaining with contrastive phrasal highlighting: A case study in assisting humans to detect translation differences. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, pages 11220–11237, Singapore. Association for Computational Linguistics.
@@
  • Carpuat et al. (2025) Marine Carpuat, Omri Asscher, Kalika Bali, Luisa Bentivogli, Frédéric Blain, Lynne Bowker, Monojit Choudhury, Hal Daumé III, Kevin Duh, Ge Gao, Alvin Grissom II, Marzena Karpinska, Elaine C. Khoong, William D. Lewis, André F. T. Martins, Mary Nurminen, Douglas W. Oard, Maja Popovic, Michel Simard, and François Yvon. 2025. An Interdisciplinary Approach to Human-Centered Machine Translation. In Proceedings of the 2025 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics.
@@
  • Pitman (2021) Jeff Pitman. 2021. Google translate: One billion installs, one billion stories. Engineering Manager, Google Translate.
@@
  • Xiao et al. (2025a) Yimin Xiao, Cartor Hancock, Sweta Agrawal, Nikita Mehandru, Niloufar Salehi, Marine Carpuat, and Ge Gao. 2025a. Sustaining human agency, attending to its cost: An investigation into generative ai design for non-native speakers’ language use. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems, CHI ’25, New York, NY, USA. Association for Computing Machinery.
  • Xiao et al. (2025b) Yimin Xiao, Yongle Zhang, Dayeon Ki, Calvin Bao, Marianna J. Martindale, Charlotte Vaughn, Ge Gao, and Marine Carpuat. 2025b. Beyond Benchmarks: Exploring Machine Translation Error Perception and Reliance Among the General Public. In Proceedings of the 2025 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics.
@@ C.1 Annotation Interface
Post-Task Survey (Control).
In the post-task survey for the control condition, we include the following two questions:
  • Mental Workload: How much mental burden (e.g., thinking, deciding) did you experience while doing the task?
    • $\circ$
      1 (Low)
    • $\circ$
    • $\circ$
      7 (High)
  • Helpfulness: How helpful was seeing the same example twice in your Spanish translation quality assessment?
    • $\circ$
      1 (Very unhelpful)
    • $\circ$
      2 (Unhelpful)
    • $\circ$
      3 (Neutral)
    • $\circ$
      4 (Helpful)
    • $\circ$
      5 (Very helpful)
Post-Task Survey (Treatment).
In the post-task survey for treatment conditions, we dynamically replace feedback with the participant’s assigned quality feedback type. The survey includes the following three questions:
@@ C.2 Recruitment Process
We conducted our user studies on the Prolific platform. For the pre-screening survey, we used the same language proficiency questions as in the pre-task survey (Appendix C) to reliably recruit monolingual English speakers. To ensure high-quality responses, we limited participation to users with a Prolific approval rate above 90% and at least 10 prior submissions. We recruited 205 participants and invited 123 who selected ‘‘I can read long and difficult texts in English without help’’ for English proficiency, and either ‘‘I cannot understand any Spanish words or sentences at all’’ or ‘‘I can read some Spanish words and very simple sentences, such as greetings and common expressions’’ for Spanish proficiency. Each pre-screening participant received 0.20 USD, totaling 55.20 USD including Prolific platform fees.
@@
Of the 123 invited participants, 91 completed the main task. Each received a base payment of 5 USD for 20 minutes of participation, with an additional 2 USD performance-based bonus for those achieving over 70% overall decision accuracy. A total of 30 participants qualified for this bonus. One participant who failed both attention check questions was compensated but excluded from the final analysis, resulting in 90 valid responses. The median task completion time was 14 minutes, corresponding to an effective pay rate of 20 USD per hour. The total cost for the main task, including Prolific fees, was 588.01 USD.
@@
Приложение F Detailed Results
F.1 Independent vs. AI-Assisted
@@ F.2 Switch Percentage
Таблица 3: Categorization of decision transitions from Independent (rows) to AI-Assisted (columns). Appropriate: Appropriate reliance; Over: Over-reliance; Under: Under-reliance.
@@
Таблица 4: Detailed switch percentage results by quality feedback type and reliance (appropriate, over-, under-reliance).
F.3 Per Shareability Label
@@ F.4 Participant Responses
Таблица 5: Detailed results for decision accuracy and CWA by condition. Ind.: Independent decision-making step; AI: AI-assisted decision-making step. Values represent mean scores with corresponding 95% confidence intervals. Best scores for each column is bold.
@@
Таблица 6: Detailed results for decision accuracy and CWA by quality feedback type and shareability label. [Uncaptioned image]: Safe to share as-is; [Uncaptioned image]: Needs bilingual review before sharing. Values represent mean scores with corresponding 95% confidence intervals. Best scores for each column is bold.
@@
Таблица 7: Participants’ responses on how they used quality feedback in their decision-making process, specifically why the feedback was perceived as confusing or unhelpful ([Uncaptioned image]) or why it was considered helpful ([Uncaptioned image]).